Insert parentheses to make each statment true:
-8 + 3 - 2 + 7 = -2
2- 3 - 4 + 1 = 4
4 - 5 + 2 - 6 - 11 = 6
-8 + 3 - 2 + 7 = -16
2 - 3 - 4 + 1 = -6
4 - 5 + 2 - 6 - 11 = 2

Answer by unlockmath(1122)

Hello,
I'll help you with the first three, then you try to do the last three. Deal?
-(8 + 3 - 2) + 7 = -2
2- (3 - 4) + 1 = 4
4 - 5 + 2 - (6 - 11) = 6
-8 + 3 - 2 + 7 = -16
2 - 3 - 4 + 1 = -6
4 - 5 + 2 - 6 - 11 = 2
Good luck!
RJ Toftness
www.math-unlock.com